OpenStudy (anonymous):

The results of a survey of 100 incoming students show the following results: 50 students own a laptop. 46 students own a tablet. 48 students own a smartphone. 22 students own at least a laptop and a tablet. 16 students own at least a tablet and a smartphone. 18 students own at least a smartphone and a laptop. 10 own all three (laptop, a tablet and a smartphone). How many students don't own any electronic devices?

OpenStudy (anonymous):

Can someone help me out with this?

OpenStudy (anonymous):

I know this problem seem hard but I need help to solve this

OpenStudy (fibonaccichick666):

nah, it's not bad. So, first we want to make a chart ok?

OpenStudy (anonymous):

Alright

OpenStudy (fibonaccichick666):

like this

OpenStudy (fibonaccichick666):

can you fill in the info we were given

OpenStudy (anonymous):

maybe

OpenStudy (fibonaccichick666):

how many students have all 3?

OpenStudy (anonymous):

My teacher told me I had to make a venn diagram

OpenStudy (fibonaccichick666):

yea you could, not my weapon of choice, but you could

OpenStudy (anonymous):

Can someone help me solve this?

OpenStudy (the_fizicx99):

@kewlgeek555 is good at graphing ^_^

OpenStudy (anonymous):

what u mean

OpenStudy (the_fizicx99):

He can help you :P you just have to use a Venn diagram, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Do you know how to solve this Kewlgeek555

OpenStudy (kewlgeek555):

Thanks for recommending me @tHe_FiZiCx99 Anyways... This is basically a math problem that likes to play with your mind, but it shall not! Just write everything that you read down like this: Total people surveyed: 100  So now we sort of mapped this out, but we are not done yet! W ehave to match the totals. Total Laptop Owners: 50 18 (ppl who own laptop & smartphone) + 22 (ppl who own laptop & tablet) + 10 (ppl who own all 3) = 50 total laptops. Okay. Then no people have to be added to the laptop thing. Why are we doing this? Because it said "AT LEAST" - meaning that it might not have given all of the owners. c; Now, this is where the assignment is wrong because if you add all the people who own a tablet, it comes to 48, but the total people who own a tablet is 46. Hm. As for the smartphones... 18 + 16 + 10 = 44, where there are 48 total ppl who own smartphones, so we are missing 4. Add the 4 and it'll be true. Now, out total is...70. 100 - 70 = 30.

OpenStudy (anonymous):

ok

OpenStudy (kewlgeek555):

Sorry if it is very confusing. >v<

OpenStudy (anonymous):

The answer I got is 52